OpenStudy (anonymous):

find the equation of straight line which passes through the points(3,-5) (-3,5)

OpenStudy (anonymous):

let's put the equation: y=ax+b, we should find a and b, the 1st point ( 3,-5) which means : -5=a(3)+b 3a+b=-5.........(1) the 2nd point is: (-3,5) which means: 5=a(-3)+b -3a+b=5.........(2) from (1): b=-5-3a.....(3) we replace it in (2): -3a+(-5-3a)=5 -6a=10 thus: a=-10/6=-5/3 we replace in (3) b= 0 so, the equation is/ \[y=(-5/3)x\]

OpenStudy (anonymous):

hey thanks ...